# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ F.1 Other Allocation Rules
In this subsection, we analyze two alternative functional forms to the proportional allocation rule, which we have extensively examined throughout the paper. We demonstrate that not only the shear proportion of allocated revenue, but also how revenue is distributed among creators, significantly influences creators’ behavior. Despite their apparent appeal, some allocation rules may not induce FSE even when the majority, if not all, of the GenAI-driven revenue is shared with creators.
Consider two alternative allocation rules parameterized by revenue share $\rho\in[0,1]$:
@@
    Winner-Takes-All (WTA): The $\rho$-fraction of AI revenue accrues to creators providing maximal effective contribution. Let $W(\mathbf{x},\mathbf{s})=\operatorname*{arg\,max}_{j\in[n]}\{x_{j}s_{j}\}$. Then:
    $f_{i,\rho}^{WTA}(\mathbf{x},\mathbf{s})=\begin{cases}\nicefrac{{\rho}}{{|W(\mathbf{x},\mathbf{s})|}}&\text{if }i\in W(\mathbf{x},\mathbf{s})\text{ and }\max_{j}x_{j}s_{j}>0\\ 0&\text{otherwise}\end{cases}$
@@
    Binary Threshold with Equal Shares (BTES): The $\rho$-fraction is shared equally among creators who fully opt-in ($s_{i}=1$). Let $S_{1}(\mathbf{s})=\{j:s_{j}=1\}$. Then:
    $f_{i,\rho}^{BTES}(\mathbf{x},\mathbf{s})=\begin{cases}\nicefrac{{\rho}}{{S_{1}(\mathbf{s})|}}&\text{if }i\in S_{1}(\mathbf{s})\\ 0&\text{otherwise}\end{cases}$
While our main results demonstrate that the proportional allocation rule admits FSEs for sufficiently high values of $\rho$ (Proposition 4, Theorem 5), the following proposition shows that such stability guarantees do not extend to these alternative allocation rules.
@@
Part 1: WTA non-existence
Consider any profile $(\mathbf{x},\mathbf{1})$ with full sharing. If all qualities are identical, any small perturbation breaks the tie and guarantees the deviator a large utility increase, so the original profile cannot have been an FSE. If qualities differ, let $i$ be a creator with $x_{i}<\max_{k}x_{k}$. Then $i\notin W(\mathbf{x},\mathbf{1})$ and receives no AI revenue. Creator $i$’s utility is:
@@
Consider deviation to $s_{i}^{\prime}=0$. The AI quality decreases to $Q_{\text{AI}}^{\prime}=\alpha(\mathbf{x}^{\top}\mathbf{s}-x_{i})<Q_{\text{AI}}$. Creator $i$ remains outside the winner set but gains from reduced AI competition:
@@
Since $Q_{\text{AI}}^{\prime}<Q_{\text{AI}}$, the fraction $\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}^{\prime}}$ increases, making deviation profitable. Thus, no full-sharing profile can be an FSE.
@@
Consider $n=2$ with costs $c_{1}(x)=0.1x^{2}$, $c_{2}(x)=0.4x^{2}$, and parameters $\alpha=\gamma=1$, $\mu=10$.
@@
First, let us derive the utility function and find the ESE $\mathbf{x}^{\star}$ (assuming $s_{1}=s_{2}=1$). In this state, $S_{1}=\{1,2\}$, so $|S_{1}|=2$. The AI quality is $Q_{\text{AI}}=\alpha(x_{1}s_{1}+x_{2}s_{2})=1(x_{1}\cdot 1+x_{2}\cdot 1)=\left\|\mathbf{x}\right\|$. The general utility for creator $i$ is:
$U_{H_{i}}(\mathbf{x},\mathbf{1};f^{BTES})=\mu(\left\|\mathbf{x}\right\|)^{\gamma}\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}}+\frac{1}{|S_{1}|}\left(\mu(\left\|\mathbf{x}\right\|)^{\gamma}\frac{Q_{\text{AI}}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}}\right)-c_{i}(x_{i}).$
Substituting the specific parameters ($\mu=10,\gamma=1,Q_{\text{AI}}=\left\|\mathbf{x}\right\|,|S_{1}|=2$):
$U_{H_{i}}(x_{1},x_{2})$ $=10(\left\|\mathbf{x}\right\|)\frac{x_{i}}{\left\|\mathbf{x}\right\|+\left\|\mathbf{x}\right\|}+\frac{1}{2}\left(10(\left\|\mathbf{x}\right\|)\frac{\left\|\mathbf{x}\right\|}{\left\|\mathbf{x}\right\|+\left\|\mathbf{x}\right\|}\right)-a_{i}x_{i}^{2}$
$=10\left\|\mathbf{x}\right\|\frac{x_{i}}{2\left\|\mathbf{x}\right\|}+\frac{1}{2}\left(10\left\|\mathbf{x}\right\|\frac{\left\|\mathbf{x}\right\|}{2\left\|\mathbf{x}\right\|}\right)-a_{i}x_{i}^{2}$
$=5x_{i}+\frac{1}{2}(5\left\|\mathbf{x}\right\|)-a_{i}x_{i}^{2}$
$=5x_{i}+2.5(x_{1}+x_{2})-a_{i}x_{i}^{2}.$
To find the qualities $(x_{1}^{\star},x_{2}^{\star})$, we solve the first-order conditions:
$\frac{\partial U_{H_{1}}}{\partial x_{1}}=5+2.5-2a_{1}x_{1}=7.5-2(0.1)x_{1}=0\implies x_{1}^{\star}=37.5,$
$\frac{\partial U_{H_{2}}}{\partial x_{2}}=5+2.5-2a_{2}x_{2}=7.5-2(0.4)x_{2}=0\implies x_{2}^{\star}=9.375.$
@@
Thus, $Q_{\text{AI}}=\left\|\mathbf{x}\right\|^{\star}=37.5+9.375=46.875$, and creator 1’s utility is:
$U_{1}(\hat{\mathbf{x}},\mathbf{1})=5(37.5)+2.5(46.875)-0.1(37.5)^{2}=164.0625.$
If creator 1 deviates to $s_{1}^{\prime}=0$, they forfeit AI revenue but the AI quality drops to $Q_{\text{AI}}^{\prime}=9.375$:
$U_{1}((37.5,9.375),(0,1))=\frac{10(46.875)(37.5)}{46.875+9.375}-0.1(37.5)^{2}=171.875.$
@@
Since $171.875>164.0625$, the deviation is profitable. Therefore, the unique ESE does not constitute an FSE under strategic sharing decisions, establishing non-existence. ∎
The failure of these allocation rules stems from misalignment between the creator’s contribution to the AI and their share of AI revenue. In the case of WTA, all the creators except one will have no incentive to contribute. The more interesting case is BTES, which compensates every contributor, but with the same shares. Thus, the strongest creators will get the same proportion of the AI revenue as the weaker ones, making them less incentivized to contribute as they can produce more content alone. Notice that the proportional allocation rule indirectly solves this misalignment, as every creator is awarded a share that is proportional to their contribution. Thus, stronger creators, who contribute more, will receive a larger share of the AI revenue.
F.2 General GenAI Quality Function
In this subsection, we present a more general formulation of the generative AI quality function and demonstrate how some of our results can be extended to this setting.
The new formulation introduces a parameter $\beta\in[0,1]$ that captures the diminishing returns of additional data. The generalized quality function is given by
$Q_{\text{AI}}=\alpha\left(\mathbf{x}^{\top}\mathbf{s}\right)^{\beta}.$
Notice that for $\beta=1$, this modeling reduces to the quality function used in the main text. In this new form, the quality growth with the shared data is no longer assumed to be linear. Most of our results cannot be directly applied to this setting. However, we present the results that somewhat do carry over.
The following is a weaker version of Proposition 2. Although it does not equip us with a full characterization of the best sharing responses of the creators, it does provide the same Prisoner’s Dilemma-like intuition discussed in the main text - profiles where no creator shares data are unstable even for not-so-large values of $\rho$.
Proposition 31.
Let $(\mathbf{x},\mathbf{0})$ be strategy profile such that $x_{i}>0$ for every $i\in[n]$. If $\rho>\frac{x_{i}}{\left\|\mathbf{x}\right\|}$ for at least one creator $i\in[n]$, then $(\mathbf{x},\mathbf{0})$ is not a PNE.
Proof of Proposition 31.
We start by characterizing the utility of an arbitrary creator $i$ within a strategy profile with no sharing $(\mathbf{x},\mathbf{0})$. Since no content is shared, the AI quality in this profile is zero. As a result, the total content quality pool simplifies to $\left\|\mathbf{x}\right\|$. Since $Q_{AI}=0$, no revenue is generated by the AI, and thus the allocated revenue is zero for all creators. Therefore, in this profile, creator $i$’s utility stems exclusively from the direct traffic attracted by their content:
$U_{i}(\mathbf{x},\mathbf{s})=T(\mathbf{x})\cdot\frac{x_{i}}{\left\|\mathbf{x}\right\|}-c_{i}(x_{i}).$
Next, we consider an unilateral deviation by creator $i$ to choosing a positive sharing level, $s_{i}^{\ast}>0$, while maintaining their quality $x_{i}$, and all the strategies of the other creators remain fixed ($s_{j}^{\prime}=0$ for $j\neq i$). Let the profile induced by this deviation be denoted as $(\mathbf{x},\mathbf{s}^{\ast})$ where $\mathbf{s}^{\ast}=(0,\ldots,s_{k}^{\ast},\ldots,0)$. In this scenario, the AI quality becomes $Q_{AI}(\mathbf{x},\mathbf{s}^{\ast})=\alpha(\sum x_{j}\cdot s_{j}^{\ast})^{\beta}=\alpha(x_{i}s_{i}^{\ast})^{\beta}$, which we represent as $z(s_{i}^{\ast})$. It is clear that $z(s_{i}^{\ast})>0$ for any $s_{i}^{\ast}>0$. The denominator correspondingly expands to $D(\mathbf{x},\mathbf{s}^{\ast})=\left\|\mathbf{x}\right\|+z(s_{i}^{\ast})$. Importantly, since the quality profile $\mathbf{x}$ remains unchanged, the total platform traffic $T=\mu(\left\|\mathbf{x}\right\|)^{\gamma}$ is unaffected by creator $i$’s deviation in sharing level $\left\|\mathbf{x}\right\|$.
Since creator $i$ is the only creator that shares its data, it receives the entire $\rho$-portion of the GenAI-driven revenue. The utility for creator $i$ under this deviation is given by:
$U_{i}(\mathbf{x},\mathbf{s}^{\ast})$ $=T\cdot\frac{x_{i}}{D(\mathbf{x},\mathbf{s}^{\ast})}+\left[T\cdot\frac{Q_{AI}(\mathbf{x},\mathbf{s}^{\ast})}{D(\mathbf{x},\mathbf{s}^{\ast})}\right]\cdot f_{k}(\mathbf{x},\mathbf{s}^{\ast})-c_{i}(x_{i})$
$=T\cdot\frac{x_{i}}{\left\|\mathbf{x}\right\|+z(s_{i}^{\ast})}+T\cdot\frac{z(s_{i}^{\ast})}{\left\|\mathbf{x}\right\|+z(s_{i}^{\ast})}\cdot\rho-c_{i}(x_{i})$
$=T\left[\frac{x_{i}+\rho z(s_{i}^{\ast})}{\left\|\mathbf{x}\right\|+z(s_{i}^{\ast})}\right]-c_{i}(x_{i}).$
To check whether this deviation is profitable for creator $i$, we compare $U_{i}(\mathbf{x},\mathbf{s}^{\ast})$ with the previous utility $U_{i}(\mathbf{x},\mathbf{s})$. Observe that the cost term $c_{i}(x_{i})$ remains constant. Thus, profitability depends on whether the revenue component has increased. Let us define the revenue function in terms of AI quality $z$ as $h(z)=T\left[\frac{x_{i}+\rho z}{\left\|\mathbf{x}\right\|+z}\right]$. We are interested in the sign of $U_{i}(\mathbf{x},\mathbf{s}^{\ast})-U_{i}(\mathbf{x},\mathbf{s})=h(z(s_{i}^{\ast}))-h(0)$.
We proceed by computing the derivative $h^{\prime}(z)$:
$h^{\prime}(z)=\frac{d}{dz}\left(T\left[\frac{x_{i}+\rho z}{\left\|\mathbf{x}\right\|+z}\right]\right)=T\cdot\frac{\rho(\left\|\mathbf{x}\right\|+z)-(x_{i}+\rho z)(1)}{(\left\|\mathbf{x}\right\|+z)^{2}}=T\cdot\frac{\rho\left\|\mathbf{x}\right\|-x_{i}}{(\left\|\mathbf{x}\right\|+z)^{2}}.$
A strictly profitable deviation for creator $i$ exists if the utility increases for some $s_{i}^{\ast}>0$. Since $h(z)$ is differentiable, this occurs if the function is increasing at $z=0$, i.e., if $h^{\prime}(0)>0$. Evaluating the derivative at $z=0$, we obtain:
$h^{\prime}(0)=T\cdot\frac{\rho\left\|\mathbf{x}\right\|-x_{i}}{(\left\|\mathbf{x}\right\|)^{2}}.$
Given that $T>0$ and $(\left\|\mathbf{x}\right\|)^{2}>0$ (since $x_{i}>0$), the sign of $h^{\prime}(0)$ is determined by the sign of its numerator. Therefore, a profitable deviation exists if $\rho\left\|\mathbf{x}\right\|-x_{i}>0$, which rearranges to the condition:
$\rho>\frac{x_{i}}{\sum_{j=1}^{n}x_{j}}.$
Another result that can be extended is Proposition 4. Turns out that it can be directly extended to this setting without the need for additional assumptions or relaxations. The main technical difference is in proving that full sharing is still the best response for all creators, as Proposition 2 cannot be applied in this setting. After establishing that, it turns out that under full sharing, the players’ utility simplifies to a function that does not depend on $\beta$ anymore, and we can employ Proposition 3 to guarantee the existence of a PNE.
Lemma 32.
Assume $\rho=1$ and $\alpha,\beta>0$. Fix a creator $i$ with $x_{i}>0$, given $\left\|\mathbf{x}_{-i}\right\|>0$. Then, $s_{i}=1$ is the unique best sharing choice.
Proof of Lemma 32.
Fix $\rho=1$. We analyze the utility function $U_{i}(s_{i})$ for creator $H_{i}$ as a function of their sharing level $s_{i}$. The utility function, excluding the constant cost term $c_{i}(x_{i})$ which is invariant w.r.t $s_{i}$, can be expressed as:
$\widehat{U}_{i}(s_{i})$ $=T(\mathbf{x})^{\gamma}\left[\frac{x_{i}}{\left\|\mathbf{x}\right\|+\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})^{\beta}}+\frac{\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}}{\left\|\mathbf{x}\right\|+\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})^{\beta}}\cdot\frac{x_{i}s_{i}}{\mathbf{x}^{\top}\mathbf{s}}\right]$
$=T(\mathbf{x})^{\gamma}\left[\frac{x_{i}}{\left\|\mathbf{x}\right\|+\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})^{\beta}}+\frac{\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta-1}x_{i}s_{i}}{\left\|\mathbf{x}\right\|+\alpha(x_{i}s_{i}+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i})^{\beta}}\right].$
Denote $T=T(\mathbf{x})$ (independent of $s_{i}$). The derivative $\frac{dU_{i}}{ds_{i}}$ is:
$\frac{dU_{i}}{ds_{i}}$ $=T\cdot\frac{N^{\prime}(s_{i})D(s_{i})-N(s_{i})D^{\prime}(s_{i})}{D(s_{i})^{2}}$
$=T\cdot\frac{\left(\alpha x_{i}\left[(\mathbf{x}^{\top}\mathbf{s})^{\beta-1}+(\beta-1)(\mathbf{x}^{\top}\mathbf{s})^{\beta-2}x_{i}s_{i}\right]\right)\left((x_{i}+\left\|\mathbf{x}_{-i}\right\|)+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}\right)}{\left(\left\|\mathbf{x}\right\|+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}\right)^{2}}$
$-T\cdot\frac{\left(x_{i}+\alpha x_{i}s_{i}(\mathbf{x}^{\top}\mathbf{s})^{\beta-1}\right)\left(\alpha\beta x_{i}(\mathbf{x}^{\top}\mathbf{s})^{\beta-1}\right)}{\left(\left\|\mathbf{x}\right\|+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}\right)^{2}}.$
Expanding the multiplications in the numerators and rearranging terms, we find that the derivative simplifies as follows:
$\frac{dU_{i}}{ds_{i}}=T\cdot\frac{\alpha x_{i}(\mathbf{x}^{\top}\mathbf{s})^{\beta-2}}{D(s_{i})^{2}}\left\{\beta x_{i}s_{i}\left\|\mathbf{x}_{-i}\right\|+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}[(1-\beta)x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}]\right\}.$
Under our assumptions ($\alpha>0$, $x_{i}>0$, $\left\|\mathbf{x}_{-i}\right\|>0$, $n\geq 2$), we have that $T\frac{\alpha x_{i}(\mathbf{x}^{\top}\mathbf{s})^{\beta-2}}{D(s_{i})^{2}}>0$ for all $s_{i}\in(0,1)$. Additionally, we have that the first expression in the curly braces is strictly positive, and the second expression is non-negative. Therefore, $\frac{dU_{i}}{ds_{i}}>0$ for all $s_{i}\in(0,1)$. Consequently, the utility function $U_{i}(s_{i})$ is strictly increasing on the interval $[0,1]$, and the unique maximum must occur at the upper boundary, $s_{i}=1$. ∎
Proof of Proposition 4.
The proof contains two parts: First, we demonstrate that a unique FSE exists under those settings, and then we show that no other PNE exists.
Existence and uniqueness of FSE.
We rely on a simplification that occurs under $\rho=1$ and full data sharing, and leverages the structure established in Proposition 3. First, we show how the utility function simplifies under full sharing. The general utility function for creator $i$ is:
$U_{i}=T(\mathbf{x})\cdot\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}}+\left[T(\mathbf{x})^{\gamma}\cdot\frac{Q_{\text{AI}}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}}\right]\cdot f_{i}(\mathbf{x},\mathbf{s})-c_{i}(x_{i}).$
Assuming full sharing, we have $\mathbf{x}^{\top}\mathbf{s}=\left\|\mathbf{x}\right\|$, and $f_{i,1}(\mathbf{x},\mathbf{s})=\frac{x_{i}}{\left\|\mathbf{x}\right\|}$. Substituting these into the general utility function:
$U_{i}(\mathbf{x})|_{\rho=1,\left\|\mathbf{x}\right\|=1}=\mu\left\|\mathbf{x}\right\|^{\gamma}\cdot\frac{x_{i}}{\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta}}+\left[\mu\left\|\mathbf{x}\right\|^{\gamma}\cdot\frac{\alpha\left\|\mathbf{x}\right\|^{\beta}}{\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta}}\right]\cdot\left(\frac{x_{i}}{\left\|\mathbf{x}\right\|}\right)-c_{i}(x_{i}).$
Combining the first two terms over a common denominator $(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})\left\|\mathbf{x}\right\|$:
$U_{i}(\mathbf{x})|_{\rho=1,\left\|\mathbf{x}\right\|=1}=\frac{\mu\left\|\mathbf{x}\right\|^{\gamma}x_{i}\left\|\mathbf{x}\right\|+\mu\left\|\mathbf{x}\right\|^{\gamma}\alpha\left\|\mathbf{x}\right\|^{\beta}x_{i}}{(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})\left\|\mathbf{x}\right\|}-c_{i}(x_{i}).$
Factoring out $\mu\left\|\mathbf{x}\right\|^{\gamma}x_{i}$ from the numerator:
$U_{i}(\mathbf{x})|_{\rho=1,\left\|\mathbf{x}\right\|=1}=\frac{\mu\left\|\mathbf{x}\right\|^{\gamma}x_{i}(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})}{(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})\left\|\mathbf{x}\right\|}-c_{i}(x_{i}).$
Given that $\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta}>0$, we can cancel the common factor $(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})$:
$U_{i}(\mathbf{x})|_{\rho=1,\left\|\mathbf{x}\right\|=1}=\frac{\mu\left\|\mathbf{x}\right\|^{\gamma}x_{i}}{\left\|\mathbf{x}\right\|}-c_{i}(x_{i})=\mu\left\|\mathbf{x}\right\|^{\gamma-1}x_{i}-c_{i}(x_{i}).$
This resulting utility is identical to the one analyzed in the proof of Proposition 3, which established uniqueness for the case where $\beta=1$. The analysis within that proof, once the utility was simplified to the form $\text{Constant}\cdot\left\|\mathbf{x}\right\|^{\gamma-1}x_{i}-c_{i}(x_{i})$, did not further depend on $\beta$. The constant factor in Proposition 6 was $K=\frac{\mu(1+\alpha\rho)}{1+\alpha}$; here, with $\rho=1$, the constant factor is simply $\mu$. Since $\mu>0$, this positive constant factor does not alter the subsequent mathematical arguments regarding concavity or uniqueness. Thus, a unique ESE exists in this setting; we denote it as $\mathbf{x}^{\star}$ for the remainder of the proof.
To show that $(\mathbf{x}^{\star},\mathbf{1})$ is the unique FSE, we employ Lemma 32. Consider an arbitrary deviation of creator $i$ to a strategy $(x^{\prime}_{i},s^{\prime}_{i})$. By a straightforward adaptation of Lemma 12 to the new form of $Q_{\text{AI}}$, we get that $\mathbf{x}^{\star}_{j}>0$ for all $j$. Thus, Lemma 32 asserts that for the deviation $(x^{\prime}_{i},s^{\prime}_{i})$ to be profitable, it must have $s^{\prime}_{i}=1$. However, notice that when all creators choose to fully share their data, the above analysis of the utility function suggests that creator $i$’s best response in choosing quality $x^{\prime}_{i}$ is to set $x^{\prime}_{i}=\mathbf{x}^{\star}_{i}$ (as $\mathbf{x}^{\star}$ is an ESE). Thus, we get that $(\mathbf{x}^{\star},\mathbf{1})$ is an FSE, and the uniqueness of the FSE stems from the uniqueness of the ESE.
No other PNE exists.
We now show that no PNE other than the FSE $(\mathbf{x}^{\star},\mathbf{1})$ can exist. Consider an arbitrary PNE $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$. By a straightforward adaptation of Lemma 12 to the new form of $Q_{\text{AI}}$, in any PNE, all creators must choose positive quality levels, so $x_{i}^{\prime}>0$ for all $i\in[n]$. Given $x_{i}^{\prime}>0$ for all $i$, $\rho=1$, and assuming $n\geq 2$ (so $\left\|\mathbf{x}_{-i}\right\|^{\prime}>0$ for each player $i$), Lemma 32 states that the unique best response for each creator $i$ regarding their sharing decision is $s_{i}^{\prime}=1$.
Therefore, any PNE must have $\mathbf{s}^{\prime}=\mathbf{1}$. As shown above, if $\mathbf{s}^{\prime}=\mathbf{1}$, the only quality profile that can sustain a PNE is the unique ESE $\mathbf{x}^{\star}$. Thus, $(\mathbf{x}^{\star},\mathbf{1})$ is the only PNE of the game. This concludes the proof of Proposition 4. ∎
Appendix G Modeling Assumptions and Limitations

### v2

@@ Appendix E Other Allocation Rules
In this section, we analyze two alternative functional forms to the proportional allocation rule examined in the main text. We demonstrate that the stability of FSEs depends not only on the total amount of revenue shared but also on how that revenue is distributed among creators. Specifically, we show that some intuitively appealing rules may fail to induce FSE even when the platform redistributes its entire GenAI-driven revenue ($\rho=1$).
We consider two alternative allocation rules parameterized by $\rho\in[0,1]$:
@@
    Winner-Takes-All (WTA): The entire $\rho$-fraction of AI revenue is allocated to the creators who provide the maximal effective contribution. Let $W(\mathbf{x},\mathbf{s})=\operatorname*{arg\,max}_{j\in[n]}\{x_{j}s_{j}\}$. Then:
    $f_{i,\rho}^{WTA}(\mathbf{x},\mathbf{s})=\begin{cases}\frac{\rho}{|W(\mathbf{x},\mathbf{s})|}&\text{if }i\in W(\mathbf{x},\mathbf{s})\text{ and }\max_{j}x_{j}s_{j}>0\\ 0&\text{otherwise}\end{cases}$
@@
    Binary Threshold with Equal Shares (BTES): The $\rho$-fraction is distributed equally among all creators who fully opt-in ($s_{i}=1$), regardless of their quality contribution. Let $S_{1}(\mathbf{s})=\{j:s_{j}=1\}$. Then:
    $f_{i,\rho}^{BTES}(\mathbf{x},\mathbf{s})=\begin{cases}\frac{\rho}{|S_{1}(\mathbf{s})|}&\text{if }i\in S_{1}(\mathbf{s})\\ 0&\text{otherwise}\end{cases}$
While Proposition˜4 and Theorem˜5 establish that the proportional allocation rule admits FSE for sufficiently high $\rho$, the following proposition shows that such stability guarantees do not extend to these alternative mechanisms.
@@
Part 1: WTA non-existence.
Consider any candidate full-sharing profile $(\mathbf{x},\mathbf{1})$. If all qualities are identical, any small perturbation breaks the tie and guarantees the deviator a large utility increase, so the original profile cannot have been an FSE. If creators have different quality levels, let $i$ be a creator such that $x_{i}<\max_{k}x_{k}$. Creator $i$ is not in the winner set $W(\mathbf{x},\mathbf{1})$ and thus receives zero AI revenue. Their utility is purely derived from direct traffic:
@@
Consider a deviation to $s_{i}^{\prime}=0$. The total shared content decreases, reducing the AI quality to $Q_{\text{AI}}^{\prime}=\alpha(\left\|\mathbf{x}\right\|-x_{i})<Q_{\text{AI}}$. Creator $i$ still receives zero AI revenue, but benefits from reduced competition:
@@
Since $Q_{\text{AI}}^{\prime}<Q_{\text{AI}}$, the traffic share $\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}^{\prime}}$ strictly increases, making the deviation profitable. If all qualities are identical, any creator can slightly increase their quality $x_{i}$ to become the unique winner, capturing the entire AI revenue pot, which also destabilizes the profile. Thus, no FSE exists.
@@
We provide a counterexample with $n=2$, $\alpha=\gamma=1$, and $\mu=10$. Let the cost functions be $c_{1}(x)=0.1x^{2}$ and $c_{2}(x)=0.4x^{2}$.
@@
First, we determine the unique ESE $\mathbf{x}^{\star}$ under full sharing ($s_{1}=s_{2}=1$). In this state, both creators share, so $|S_{1}|=2$. The AI quality is $Q_{\text{AI}}=\alpha(x_{1}+x_{2})=\left\|\mathbf{x}\right\|$. The utility for creator $i$ is:
$U_{i}(\mathbf{x},\mathbf{1};f^{BTES})=10\left\|\mathbf{x}\right\|\frac{x_{i}}{2\left\|\mathbf{x}\right\|}+\frac{1}{2}\left(10\left\|\mathbf{x}\right\|\frac{\left\|\mathbf{x}\right\|}{2\left\|\mathbf{x}\right\|}\right)-c_{i}(x_{i})=5x_{i}+2.5\left\|\mathbf{x}\right\|-c_{i}(x_{i}).$
The first-order conditions $\frac{\partial U_{i}}{\partial x_{i}}=0$ yield:
$5+2.5-0.2x_{1}=0\implies x_{1}^{\star}=37.5,$
$5+2.5-0.8x_{2}=0\implies x_{2}^{\star}=9.375.$
@@
The resulting total quality is $\left\|\mathbf{x}\right\|^{\star}=46.875$. Creator 1’s utility in this potential FSE is:
$U_{1}(\mathbf{x}^{\star},\mathbf{1})=5(37.5)+2.5(46.875)-0.1(37.5)^{2}=187.5+117.1875-140.625=164.0625.$
Now consider a deviation by Creator 1 to $s_{1}^{\prime}=0$ (keeping $x_{1}=37.5$). They forfeit their share of AI revenue, but the AI quality drops to $Q_{\text{AI}}^{\prime}=\alpha x_{2}^{\star}=9.375$. Creator 1’s new utility is derived solely from direct traffic:
$U_{1}^{dev}=10(46.875)\frac{37.5}{46.875+9.375}-0.1(37.5)^{2}=468.75\left(\frac{37.5}{56.25}\right)-140.625=171.875.$
@@
Thus, the deviation is profitable. Thus, the unique ESE is not stable under strategic sharing, proving the non-existence of FSE. ∎
The instability of these rules stems from a misalignment between contribution and reward. Under WTA, non-winning creators have no incentive to share, as doing so only strengthens the GenAI competitor without providing any compensation. Under BTES, high-quality creators (like Creator 1 in the proof) are undercompensated relative to their large contribution to the AI’s quality; they prefer to withhold data to weaken the AI rather than accept a diluted share of revenue equal to that of lower-quality contributors. The proportional allocation rule resolves this by linking rewards directly to the magnitude of contribution, ensuring that stronger creators are adequately incentivized.
Appendix F Modeling Assumptions and Limitations